OpenStudy (anonymous):

find all solutions of the equation 2cos 3x =1 in the interval [0,pi).

OpenStudy (amistre64):

when does cos(t) = 1/2 ?

OpenStudy (amistre64):

and check the interval, is that spose to be pi or 2pi

OpenStudy (anonymous):

pi/3 and 5pi/3 thas wher cos 1/2

OpenStudy (amistre64):

then 3x needs to be pi/3 or 5pi/3

OpenStudy (anonymous):

the interval is 0,pi

OpenStudy (amistre64):

then when 3x = pi/3 is our only concern

OpenStudy (amistre64):

2cos 3x =1 cos 3x =1/2 cos t =1/2, t = pi/3 3x = pi/3

OpenStudy (anonymous):

thnks